Circular No. 5830

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRS 1915+105 I. F. Mirabel, P. A. Duc, R. Teyssier, and J. Paul, Service d’Astrophy-

sique, Centre d’Etudes de Saclay; L. F. Rodriguez, Universidad Nacional Autónoma de México; J. Marti, Departament d’Astronomia i Meteorologia, University of Barcelona; D. Golombek, Space Telescope Science Institute; and M. Aurière, Observatoire du Pic du Midi, communicate: “Infrared images in J, H, and K taken with the 2.2-m telescope (+ IRAC2) at the European Southern Observatory revealed a variable infrared source at α = 19^{h}12^{m}50^{s}.01, δ = +10^{o}51′27′′.5 ± 1′′ (equinox 1950.0), close to the variable 20-cm radio source previously reported by us (IAUC 5773). The infrared and radio sources are inside the ROSAT 10′′ error box for the Aquila x-ray transient (IAUC 5786). During the night of June 3-4, we measured J > 19, H = 16.2, K = 14.3; during June 4-5, J > 19, H = 15.0, K = 13.4. This is the only infrared source in the field that showed significant day-to-day variations: 1.2 mag in H and 0.9 in K. New Very Large Array observations of the radio counterpart on Apr. 27 showed that at 20 cm it went below a detection limit of 0.25 mJy (a factor of 30 below its flux during Mar. 9-10). However, on that date at 3.6 cm, we measured a flux 0.20 mJy ± 0.02 centered at α = 19^{h}12^{m}49^{s}.98, δ = +10^{o}51′26′′.8 (± 0′′.03), which implies a spectral index > –0.2. A displacement of 1′′.2 between the position of the radio source at 20 cm in 1992 December and its position at 3.6 cm in 1993 April could be due to separate radio brightenings along jets emanating from the x-ray transient. VLA and infrared observations of this source will be made on Aug. 24.”

A. J. Castro-Tirado, Danish Space Research Institute (DSRI) and Laboratorio de Astrofisica Espacial, LAEFF, Spain; J. Davies, Joint Astromony Centre, Hilo; and S. Brandt and N. Lund, DSRI, write: “Using IRCAM images taken with the 3.6-m U.K. Infrared Telescope at Mauna Kea by C. Everall on July 7, we have found a possible counterpart for the variable radio source located at α = 19^{h}12^{m}49^{s}.97, δ = +10^{o}51′25′′.6 (equinox 1950.0), inside the x-ray error box (cf. IAUC 5786) of GRS 1915+105, discovered on 1992 Aug. 15 (IAUC 5590, 5619, 5773). The infrared counterpart (having magnitudes J = 16.61 ± 0.08, K = 12.95 ± 0.05) agrees with the position of the radio source within 0′′.5, located 2′′.3 northwest of a star of V = 20.5. On May 10-12, EURECA WATCH observations showed that GRS 1915+105 was still active in x-rays (see also IAUC 5779).”

1993 July 16 Daniel W. E. Green